OpenStudy (anonymous):

Could someone PLEASE help me? I dont understand how to do this! Sheila has taken five tests in each subject this semester. Match each subject with the number of points the mean would change if the lowest grade is dropped. (round to the nearest whole number) Match Term Definition) Math: 28, 88, 78, 90, 78 Science: 84, 75, 81, 77, 29 Language Arts: 95, 84, 89, 45, 92 History: 88, 92, 48, 85, 85 A) 10 B) 11 C) 8 D) 9

OpenStudy (anonymous):

Just think about it, read the question a couple of times until you figure it out. If you don't know how to find the average you add up the numbers and divide by however many numbers you added up

OpenStudy (anonymous):

Okay, I did the (mean) part and I got for math was 72.4, Science was 69.2 Language arts was 81 and History was 79.6, I just dont know what 10, 11, 8 and 9 are for.

OpenStudy (texaschic101):

Math -- 28 + 88 + 78 + 90 + 78 = 362. The mean is 362/5 = 72.4 If the lowest grade was dropped the mean would be ...362 - 28 = 334. Now divide 334 by 4 and you get 83.5. And the difference of the two means are .. 83.5 - 72.4 = 11.1 = 11 (answer B) Now do this to all the subjects...find the difference between the two means, and you will get your answers. Do you understand ?